# Energy Extraction via Magnetic Reconnection in Konoplya-Rezzolla-Zhidenko Parameterized Black Holes

### v1

@@ IV Energy extraction via magnetic reconnection
In this section, we investigate the process of harnessing energy from the black hole through magnetic reconnection in the ergoregion, a phenomenon anticipated to be a common occurrence in rapidly spinning black holes. To simplify the analysis, we employ the methodology elucidated in [22, 23] and focus on magnetic reconnection in the surrounding plasma, which rotates around the black hole in a stable circular orbit in the equatorial plane at the Keplerian velocity $\Omega_{K}$.
To assess the energy density of the outflowing plasma, it is advantageous to evaluate some quantities in the so-called “zero-angular-momentum-observer” (ZAMO) frame [9]. This is a locally non-rotating frame in which the spacetime is locally Minkowskian
@@
and $\omega^{\phi}=-g_{t\phi}/g_{\phi\phi}$ is the angular velocity of the frame dragging. We denote quantities in ZAMO frame with hats. The Keplerian velocity of the co-rotating bulk plasma in ZAMO becomes
@@
where $p,\omega,U^{\mu}$ and $F^{\mu\nu}$ are the proper plasma pressure, enthalpy density, four-velocity, and electromagnetic field tensors, respectively. With the time-like killing vector $\xi=\partial_{t}$, one can define a covariant conserved energy current $J^{\mu}\equiv T^{\mu\nu}\xi_{\nu}$ and the associated energy density $e^{\infty}\equiv n_{\mu}J^{\mu}=-\alpha g_{\mu 0}T^{\mu 0}$, where $n^{\mu}$ is the unit vector normal to time-like hypersurfaces $t=consant$. At infinity, $e^{\infty}$ is the energy density measured by static observers, and so is usually called the “energy-at-infinity” density. Assuming a highly efficient conversion of magnetic energy into kinetic energy, so that the electromagnetic component of the energy density can be neglected, and taking the “relativistic adiabatic incompressible sphere approach”, the energy-at-infinity density can be expressed as [22, 23]
@@
where $\xi$ is the orientation angle between the outflow plasma velocity and the azimuthal direction in the equatorial plane in the local rest frame, $\hat{\gamma}_{K}=(1-\hat{v}_{K}^{2})^{-1/2}$ and $\sigma_{0}$ is the plasma magnetization. The signs $+$ and -- stand for accelerated and decelerated parts, respectively. From the above equation, it can be seen that the energy-at-infinity density is parameterized by five parameters $\{a,\delta_{1},\delta_{2},\sigma_{0},\xi\}$ as well as the radius $r=r_{X}$ of the circular orbit of the plasma where the reconnection occurs (X-point).
@@ IV.2 Energy extraction power and efficiency
The power ${\cal P}_{\rm extr}$ per unit enthalpy extracted from the black hole by the escaping plasma can be well estimated as [23]
@@
where $A_{\rm in}$ is the cross-sectional area of the inflowing plasma, which can be estimated as $A_{\rm in}\sim(r_{\rm sl}^{2}-r_{\rm ISCO}^{2})$ for fast-rotating black holes. The parameter $U_{\rm in}={\cal O}(10^{-2})$ and ${\cal O}(10^{-1})$ for the collisional and collisionless regime respectively.
Fig. 6 shows a typical picture of power ${\cal P}_{\rm extr}$ as a function of the X-point location $r_{X}$. From the figure it can be seen that, with all other parameters fixed, ${\cal P}_{\rm extr}$ is a monotonically decreasing function of $r_{X}$, reaching a maximum at $r_{X}=r^{+}_{\rm ISCO}$. From the left panels, we can see that for fixed $r_{X}$ the effect of $\delta_{1}$ on ${\cal P}_{\rm extr}$ is moderate and depends on the spin $a$. When $a$ is not so extremely high, ${\cal P}_{\rm extr}$ is a monotonically decreasing function of $\delta_{1}$. However, when $a$ is extremely high, as $\delta_{1}$ increases, ${\cal P}_{\rm extr}$ first increases and then decreases. From the right panels, we can see that for fixed $r_{X}$, ${\cal P}_{\rm extr}$ is a monotonically increasing function of $\delta_{2}$. Note that for $a<0.994$, as we mentioned above, there is no energy extraction for $\delta_{2}=-0.36$. The maximum power for the KRZ black hole can significantly exceed that of the Kerr case. For instance, when $a=0.99$, ${\cal P}^{\rm max}_{\rm extr}(\delta_{1}=-0.26,\delta_{2}=0)/{\cal P}^{\rm max}_{\rm extr}({\rm Kerr})\sim 2.0$ and ${\cal P}^{\rm max}_{\rm extr}(\delta_{1}=0,\delta_{2}=0.21)/{\cal P}^{\rm max}_{\rm extr}({\rm Kerr})\sim 1.7$. Similarly, for $a=0.995$, ${\cal P}^{\rm max}_{\rm extr}(\delta_{1}=-0.1,\delta_{2}=0)/{\cal P}^{\rm max}_{\rm extr}({\rm Kerr})\sim 1.5$ and ${\cal P}^{\rm max}_{\rm extr}(\delta_{1}=0,\delta_{2}=0.21)/{\cal P}^{\rm max}_{\rm extr}({\rm Kerr})\sim 1.5$.
To further measure the feasibility of magnetic reconnection energy extraction, it is convenient to define the efficiency of the plasma energization process [23]
@@
If $\eta>1$, energy is extracted from the black hole. Fig. 7 shows a typical picture of the efficiency $\eta$ as a function of the X-point location $r_{X}$. From the figure it can be seen that, for other fixed parameters, $\eta$ is a monotonically decreasing function of $r_{X}$, reaching a maximum at $r_{X}=r^{+}_{\rm ISCO}$. The right panels show that for fixed $r_{X}$, $\eta$ is a monotonically increasing function of $\delta_{2}$. However, the effect of $\delta_{1}$ on $\eta$ is complicated and moderately dependent on $r_{X}$. When $r_{X}$ is small, $\eta$ is a monotonically increasing function of $\delta_{1}$. However, if $r_{X}$ is large enough, $\eta$ becomes a monotonically decreasing function of $\delta_{1}$. The maximum efficiency for the KRZ black hole can also exceed that of the Kerr case. For instance, when $a=0.99$, $\eta^{\rm max}(\delta_{1}=-0.26,\delta_{2}=0)/\eta^{\rm max}({\rm Kerr})\sim 105\%$ and $\eta^{\rm max}(\delta_{1}=0,\delta_{2}=0.21)/\eta^{\rm max}({\rm Kerr})\sim 106\%$. Similarly, for $a=0.995$, $\eta^{\rm max}(\delta_{1}=-0.1,\delta_{2}=0)/\eta^{\rm max}({\rm Kerr})\sim 106\%$ and $\eta^{\rm max}(\delta_{1}=0,\delta_{2}=0.21)/\eta^{\rm max}({\rm Kerr})\sim 109\%$.
@@ V Summary and conclusions
In this work, we study energy extraction from the Konoplya-Rezzolla-Zhidenko (KRZ) parameterized black holes via the Comisso-Asenjo mechanism [23], a novel mechanism recently proposed to extract energy from black holes via the magnetic reconnection process. The KRZ metric is a parametrized metric that describes generic stationary and axisymmetric black holes [50], with six parameters introduced to characterize deviations from the Kerr metric. Of these parameters, $\delta_{1}$ and $\delta_{2}$ play a crucial role in the physics under consideration, corresponding to deformations of $g_{tt}$ and rotational deformations of the metric, respectively, especially in the near-horizon zone. The values of both parameters are constrained by current astronomical observations (8). We study in detail the ergoregion, the circular geodesics in the equatorial plane, and the energy extraction via magnetic reconnection, and find that these two deformation parameters have significant influences.
Magnetic reconnection takes place in the ergoregion. Following [22], we have assumed that the bulk plasma rotates around the black hole in a circular and stable orbit in the equatorial plane. Such an orbit exists from infinity to ISCO. As shown in Figs. 1, 2 and 3, $\{\delta_{1},\delta_{2}\}$ significantly affects the shape of the ergoregion and the location of ISCO. $\{\delta_{1},\delta_{2}\}$ can enlarge or shrink the ergoregion depending on their signs and the spin $a$. Specifically, a negative $\delta_{1}$ and a positive $\delta_{2}$ promote magnetic reconnection by enlarging the ergoregion. The counter-rotating ISCO $r^{-}_{\rm ISCO}$ is always outside the ergoregion and is therefore not relevant to the process. The co-rotating ISCO enters the ergoregion when the spin $a$ is high enough to exceed some critical value $a_{c}$. $\delta_{1}$ has little effect on $a_{c}$, while $\delta_{2}$ has a significant effect on $a_{c}$. A higher positive value of $\delta_{2}$ is advantageous for magnetic reconnection, as it necessitates a lower black hole spin for the process to occur.
@@
From Fig. 4 it can be seen that an increase in the magnitude of negative $\delta_{1}$ or positive $\delta_{2}$ leads to a reduction in the critical plasma magnetization $\sigma_{0}$ required for energy extraction. This observation indicates a preference for energy extraction with negative $\delta_{1}$ and positive $\delta_{2}$. From Fig. 5 it can be seen that as $\delta_{2}$ decreases from positive to negative, the permissible region in the $a-r_{X}$ plane to satisfy the energy extraction conditions shrinks significantly and a higher value of $a$ is required. It is noteworthy that if $\delta_{2}$ is sufficiently negative, an exceedingly high spin (e.g. $a\gtrsim 0.994$ for $\delta_{2}=-0.36$) is required to satisfy the energy extraction conditions, posing challenges for energy extraction via this mechanism. This again confirms our previous conclusion that energy extraction is favored by a large positive $\delta_{2}$. This conclusion is further confirmed by examining the effects of the deformation parameters $\{\delta_{1},\delta_{2}\}$ on the power ${\cal P}_{\rm extr}$ (Fig. 6) and efficiency $\eta$ (Fig. 7). All these results imply that energy extraction is favored by negative $\delta_{1}$ but positive $\delta_{2}$. Compared to the Kerr case, the maximum power and efficiency for the KRZ black hole can be significantly higher.
Typically, the energy of the matter and electromagnetic fields surrounding a black hole is negligible compared to the mass of the black hole. Therefore, in our work, as in the related studies mentioned earlier, we have not considered the back-reactions of the surrounding matter and electromagnetic fields on the black hole geometry. Non-Kerr rotating black holes typically have only a small deviation from the Kerr black hole. If the energy of the matter and fields surrounding these black holes is sufficiently high, their back-reaction on the geometry can be comparable to the deviation. In this case, it is necessary to consider their effects on the black hole geometry and the energy extraction process. Further studies are needed to fully understand these effects.

### v2

@@ IV Energy extraction via magnetic reconnection
In this section, we investigate the process of harnessing energy from the black hole through magnetic reconnection in the ergoregion, a phenomenon anticipated to be a common occurrence in rapidly spinning black holes. To simplify the analysis, we employ the methodology elucidated in [23] and focus on magnetic reconnection in the surrounding plasma, which rotates around the black hole in a stable circular orbit in the equatorial plane at the Keplerian velocity $\Omega_{K}$.
To assess the energy density of the outflowing plasma, it is advantageous to evaluate some quantities in the so-called “zero-angular-momentum-observer” (ZAMO) frame [9]. This is a locally nonrotating frame in which the spacetime is locally Minkowskian
@@
and $\omega^{\phi}=-g_{t\phi}/g_{\phi\phi}$ is the angular velocity of the frame dragging. We denote quantities in ZAMO frame with hats. The Keplerian velocity of the corotating bulk plasma in ZAMO becomes
@@
where $p,\omega,U^{\mu}$ and $F^{\mu\nu}$ are the proper plasma pressure, enthalpy density, four-velocity, and electromagnetic field tensors, respectively. With the timelike killing vector $\xi=\partial_{t}$, one can define a covariant conserved energy current $J^{\mu}\equiv T^{\mu\nu}\xi_{\nu}$ and the associated energy density $e^{\infty}\equiv n_{\mu}J^{\mu}=-\alpha g_{\mu 0}T^{\mu 0}$, where $n^{\mu}$ is the unit vector normal to timelike hypersurfaces $t=consant$. At infinity, $e^{\infty}$ is the energy density measured by static observers, and so is usually called the “energy-at-infinity” density. Assuming a highly efficient conversion of magnetic energy into kinetic energy, so that the electromagnetic component of the energy density can be neglected, and adopting the approximation that the plasma is incompressible and adiabatic, the energy-at-infinity density can be expressed as [23]
@@
where $\xi$ is the orientation angle between the outflow plasma velocity and the azimuthal direction in the equatorial plane in the local rest frame, $\hat{\gamma}_{K}=(1-\hat{v}_{K}^{2})^{-1/2}$ and $\sigma_{0}$ is the plasma magnetization. The signs $+$ and -- stand for accelerated and decelerated parts, respectively. From the above equation, it can be seen that the energy-at-infinity density is parametrized by five parameters $\{a,\delta_{1},\delta_{2},\sigma_{0},\xi\}$ as well as the radius $r=r_{X}$ of the circular orbit of the plasma where the reconnection occurs (X-point).
@@ IV.2 Energy extraction power and efficiency
The power ${\cal P}_{\rm extr}$ per unit enthalpy extracted from the black hole by the escaping plasma can be well-estimated as [23]
@@
where $A_{\rm in}$ is the cross sectional area of the inflowing plasma, which can be estimated as $A_{\rm in}\sim(r_{\rm sl}^{2}-r_{\rm ISCO}^{2})$ for fast-rotating black holes. The parameter $U_{\rm in}={\cal O}(10^{-2})$ and ${\cal O}(10^{-1})$ for the collisional and collisionless regime respectively.
Figure 6 shows a typical picture of power ${\cal P}_{\rm extr}$ as a function of the X-point location $r_{X}$. From the figure it can be seen that, with all other parameters fixed, ${\cal P}_{\rm extr}$ is a monotonically decreasing function of $r_{X}$, reaching a maximum at $r_{X}=r^{+}_{\rm ISCO}$. From the left panels, we can see that for fixed $r_{X}$ the effect of $\delta_{1}$ on ${\cal P}_{\rm extr}$ is moderate and depends on the spin $a$. When $a$ is not so extremely high, ${\cal P}_{\rm extr}$ is a monotonically decreasing function of $\delta_{1}$. However, when $a$ is extremely high, as $\delta_{1}$ increases, ${\cal P}_{\rm extr}$ first increases and then decreases. From the right panels, we can see that for fixed $r_{X}$, ${\cal P}_{\rm extr}$ is a monotonically increasing function of $\delta_{2}$. Note that for $a<0.994$, as we mentioned above, there is no energy extraction for $\delta_{2}=-0.36$. For given $a$, the maximum power for the KRZ black hole can significantly exceed that of the Kerr case when $\{\delta_{1},\delta_{2}\}$ fall within particular ranges. For instance, when $a=0.99$, $\frac{{\cal P}^{\rm max}_{\rm extr}(\delta_{1}=-0.26,\delta_{2}=0)}{{\cal P}^{\rm max}_{\rm extr}({\rm Kerr})}\sim 2.0$ and $\frac{{\cal P}^{\rm max}_{\rm extr}(\delta_{1}=0,\delta_{2}=0.21)}{{\cal P}^{\rm max}_{\rm extr}({\rm Kerr})}\sim 1.7$. Similarly, for $a=0.995$, $\frac{{\cal P}^{\rm max}_{\rm extr}(\delta_{1}=-0.1,\delta_{2}=0)}{{\cal P}^{\rm max}_{\rm extr}({\rm Kerr})}\sim 1.5$ and $\frac{{\cal P}^{\rm max}_{\rm extr}(\delta_{1}=0,\delta_{2}=0.21)}{{\cal P}^{\rm max}_{\rm extr}({\rm Kerr})}\sim 1.5$.
To further measure the feasibility of magnetic reconnection energy extraction, it is convenient to define the efficiency of the plasma energization process [23],
@@
If $\eta>1$, energy is extracted from the black hole. Figure 7 shows a typical picture of the efficiency $\eta$ as a function of the X-point location $r_{X}$. From the figure it can be seen that, for other fixed parameters, $\eta$ is a monotonically decreasing function of $r_{X}$, reaching a maximum at $r_{X}=r^{+}_{\rm ISCO}$. The right panels show that for fixed $r_{X}$, $\eta$ is a monotonically increasing function of $\delta_{2}$. However, the effect of $\delta_{1}$ on $\eta$ is complicated and moderately dependent on $r_{X}$. When $r_{X}$ is small, $\eta$ is a monotonically increasing function of $\delta_{1}$. However, if $r_{X}$ is large enough, $\eta$ becomes a monotonically decreasing function of $\delta_{1}$. For given $a$, the maximum efficiency for the KRZ black hole can also exceed that of the Kerr case when $\{\delta_{1},\delta_{2}\}$ fall within particular ranges. For instance, when $a=0.99$, $\frac{\eta^{\rm max}(\delta_{1}=-0.26,\delta_{2}=0)}{\eta^{\rm max}({\rm Kerr})}\sim 105\%$ and $\frac{\eta^{\rm max}(\delta_{1}=0,\delta_{2}=0.21)}{\eta^{\rm max}({\rm Kerr})}\sim 106\%$. Similarly, for $a=0.995$, $\frac{\eta^{\rm max}(\delta_{1}=-0.1,\delta_{2}=0)}{\eta^{\rm max}({\rm Kerr})}\sim 106\%$ and $\frac{\eta^{\rm max}(\delta_{1}=0,\delta_{2}=0.21)}{\eta^{\rm max}({\rm Kerr})}\sim 109\%$.
@@ V Summary and conclusions
In this work, we study energy extraction from the KRZ parametrized black holes via the Comisso-Asenjo mechanism [23], a novel mechanism recently proposed to extract energy from black holes via the magnetic reconnection process. The KRZ metric is a parametrized metric that describes generic stationary and axisymmetric black holes [50], with six parameters introduced to characterize deviations from the Kerr metric. Of these parameters, $\delta_{1}$ and $\delta_{2}$ play a crucial role in the physics under consideration, corresponding to deformations of $g_{tt}$ and rotational deformations of the metric, respectively, especially in the near-horizon zone. The values of both parameters are constrained by current astronomical observations (8). We study in detail the ergoregion, the circular geodesics in the equatorial plane, and the energy extraction via magnetic reconnection, and find that these two deformation parameters have significant influences.
Magnetic reconnection takes place in the ergoregion. Following [22], we have assumed that the bulk plasma rotates around the black hole in a circular and stable orbit in the equatorial plane. Such an orbit exists from infinity to ISCO. As shown in Figs. 1, 2 and 3, $\{\delta_{1},\delta_{2}\}$ significantly affects the shape of the ergoregion and the location of ISCO. $\{\delta_{1},\delta_{2}\}$ can enlarge or shrink the ergoregion depending on their signs and the spin $a$. Specifically, a negative $\delta_{1}$ and a positive $\delta_{2}$ promote magnetic reconnection by enlarging the ergoregion. The counter-rotating ISCO $r^{-}_{\rm ISCO}$ is always outside the ergoregion and is therefore not relevant to the process. The corotating ISCO enters the ergoregion when the spin $a$ is high enough to exceed some critical value $a_{c}$. $\delta_{1}$ has little effect on $a_{c}$, while $\delta_{2}$ has a significant effect on $a_{c}$. A higher positive value of $\delta_{2}$ is advantageous for magnetic reconnection, as it necessitates a lower black hole spin for the process to occur.
@@
From Fig. 4 it can be seen that an increase in the magnitude of negative $\delta_{1}$ or positive $\delta_{2}$ leads to a reduction in the critical plasma magnetization $\sigma_{0}$ required for energy extraction. This observation indicates a preference for energy extraction with negative $\delta_{1}$ and positive $\delta_{2}$. From Fig. 5 it can be seen that as $\delta_{2}$ decreases from positive to negative, the permissible region in the $a-r_{X}$ plane to satisfy the energy extraction conditions shrinks significantly and a higher value of $a$ is required. It is noteworthy that if $\delta_{2}$ is sufficiently negative, an exceedingly high spin (e.g., $a\gtrsim 0.994$ for $\delta_{2}=-0.36$) is required to satisfy the energy extraction conditions, posing challenges for energy extraction via this mechanism. This again confirms our previous conclusion that energy extraction is favored by a large positive $\delta_{2}$. This conclusion is further confirmed by examining the effects of the deformation parameters $\{\delta_{1},\delta_{2}\}$ on the power ${\cal P}_{\rm extr}$ (Fig. 6) and efficiency $\eta$ (Fig. 7). All these results imply that energy extraction is favored by negative $\delta_{1}$ but positive $\delta_{2}$. Compared to the Kerr case, the maximum power and efficiency for the KRZ black hole can be significantly higher.
Typically, the energy of the matter and electromagnetic fields surrounding a black hole is negligible compared to the mass of the black hole. Therefore, in our work, as in the related studies mentioned earlier, we have not considered the backreactions of the surrounding matter and electromagnetic fields on the black hole geometry. Non-Kerr rotating black holes typically have only a small deviation from the Kerr black hole. If the energy of the matter and fields surrounding these black holes is sufficiently high, their backreaction on the geometry can be comparable to the deviation. In this case, it is necessary to consider their effects on the black hole geometry and the energy extraction process. Further studies are needed to fully understand these effects.